# Stochastic Control Methods for Optimization

### v1

@@ 3.3 Numerical approximation with examples
$\mathcal{L}^{N}[w^{N}_{\varepsilon}]$ $=N\left[-\partial_{t}\mathcal{V}_{\varepsilon}-\frac{1}{2}\int\text{div}_{y}(\partial_{\mu}\mathcal{V}_{\varepsilon})d\mu^{N}+\frac{1}{2\varepsilon}\int|\partial_{\mu}\mathcal{V}_{\varepsilon}|^{2}d\mu^{N}\right]$
$\quad-\frac{1}{2N}\sum_{i=1}^{N}\text{Tr}(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon})(x_{i},x_{i}).$
Notice that the term in the square brackets $\dots$] is exactly the master equation evaluated at $\mu=\mu^{N}_{\mathbf{x}}$. Since $\mathcal{V}_{\varepsilon}$ is a solution to the master equation, this bracketed term is zero. The remaining term is the residual error:
$R^{N}_{\varepsilon}(t,\mathbf{x})=-\frac{1}{2N}\sum_{i=1}^{N}\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}},x_{i},x_{i})\right).$ (A.4)
The boundedness of $R^{N}_{\varepsilon}$ depends on the regularity of the master equation solution $\mathcal{V}_{\varepsilon}$ in Assumption 3.1. Thus, we have
$|R^{N}_{\varepsilon}(t,\mathbf{x})|\leq\frac{1}{2N}\cdot N\cdot|\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}\right)|\leq\frac{1}{2}L_{\varepsilon}.$
Therefore, the equation satisfied by $w^{N}_{\varepsilon}$ is $\mathcal{L}^{N}[w^{N}_{\varepsilon}]=R^{N}_{\varepsilon}$, where the residual $R^{N}_{\varepsilon}$ is uniformly bounded in $N$ but may depend on $\varepsilon$, while the equation satisfied by the value function $v^{N}_{\varepsilon}$ is $\mathcal{L}^{N}[v^{N}_{\varepsilon}]=0.$
@@
$\begin{cases}\partial_{t}e^{N}_{\varepsilon}+\frac{1}{2}\sum_{i=1}^{N}\Delta_{x_{i}}e^{N}_{\varepsilon}+\sum_{i=1}^{N}\mathbf{b}_{i}\cdot\partial_{x_{i}}e^{N}_{\varepsilon}=R^{N}_{\varepsilon},\\ e^{N}_{\varepsilon}(1,\mathbf{x})=0.\end{cases}$ (A.5)
@@
In this way, we have established that for the total value functions:
@@
Finally, dividing by $N$ to get the normalized estimate:

### v2

@@ 3.3 Numerical approximation with examples
$\mathcal{L}^{N}[w^{N}_{\varepsilon}](t,\mathbf{x})$ $=N\left[-\partial_{t}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})-\frac{1}{2}\int_{\mathbb{R}^{d}}\text{div}_{y}(\partial_{\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(y)\mu^{N}_{\mathbf{x}}(dy)+\frac{1}{2\varepsilon}\int_{\mathbb{R}^{d}}|\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})(y)|^{2}\mu^{N}_{\mathbf{x}}(dy)\right]$
$\qquad-\frac{1}{2}\int_{\mathbb{R}^{d}}\text{Tr}(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(y,y)\mu^{N}_{\mathbf{x}}(dy).$
Notice that the term in the square brackets $\dots$] is exactly the master equation evaluated at $\mu=\mu^{N}_{\mathbf{x}}$. Since $\mathcal{V}_{\varepsilon}$ is a solution to the master equation, this bracketed term is zero. The remaining term is the residual error
$R^{N}_{\varepsilon}(t,\mathbf{x})=-\frac{1}{2}\int_{\mathbb{R}^{d}}\text{Tr}(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon})(t,\mu^{N}_{\mathbf{x}})(y,y)\mu^{N}_{\mathbf{x}}(dy),$ (B.4)
whose boundedness depends on the regularity of the master equation solution $\mathcal{V}_{\varepsilon}$ in Assumption 3.1. Thus, we have
$|R^{N}_{\varepsilon}(t,\mathbf{x})|\leq\frac{1}{2}\left|\int_{\mathbb{R}^{d}}\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu^{N}_{\mathbf{x}})(y,y)\right)\mu^{N}_{\mathbf{x}}(dy)\right|\leq\frac{1}{2}L_{\varepsilon}.$
It is straightforward to check that the function $v^{N}_{\varepsilon}$ defined through the (inverse) Cole-Hopf transformation in (1.24) is a classical solution to the HJB equation (1.23). Therefore, the equation satisfied by $w^{N}_{\varepsilon}$ is $\mathcal{L}^{N}[w^{N}_{\varepsilon}]=R^{N}_{\varepsilon}$, where the residual $R^{N}_{\varepsilon}$ is uniformly bounded in $N$ but may depend on $\varepsilon$, while the equation satisfied by the value function $v^{N}_{\varepsilon}$ is $\mathcal{L}^{N}[v^{N}_{\varepsilon}]=0.$
@@
$\begin{cases}\partial_{t}e^{N}_{\varepsilon}+\frac{1}{2}\sum_{i=1}^{N}\Delta_{x_{i}}e^{N}_{\varepsilon}+\sum_{i=1}^{N}\mathbf{b}_{i}\cdot\partial_{x_{i}}e^{N}_{\varepsilon}=R^{N}_{\varepsilon},\\ e^{N}_{\varepsilon}(1,\mathbf{x})=0.\end{cases}$ (B.5)
@@
In this way, we have established that
@@
Finally, dividing by $N$ to reach the normalized estimate